OpenStudy (anonymous):

**GIVING MEDAL** Please help, I’m very confused :( What is the standard form of the equation of the conic given by 2x^2 + 2y^2 + 4x – 12y – 22 = 0

OpenStudy (anonymous):

Choices: (x + 1)^2/21 – (y - 3)^2/21 = 1 (x + 1)^2/21 + (y - 3)^2/21 = 1 (x - 3)^2/21 + (y + 1)^2/21 = 1 (x - 1)^2/7 + (y + 3)^2/3 = 1

OpenStudy (anonymous):

combine x term and y term seperatly

OpenStudy (anonymous):

first take 2 common

OpenStudy (anonymous):

after taking 2 commonweare left with x^2+y^2+2x-6y-11=0

OpenStudy (anonymous):

x^2+2*x*1+1-1+y^2-2*3*y+9-9-11=0 x^2+2*x*1=(x+1)^2 y^2-2*3*y+9=(y-3)^2 (x+1)^2+(y-3)^2=11+9+1=21 (x+1)^2/21+(y-3)^2/21=1

OpenStudy (anonymous):

Ah, thank you so much! I actually understand it now!